OpenStudy (melissa_something):

Complete the square and write equation in standard form and give center and radius. -___- PLEASE HELP

OpenStudy (melissa_something):

x^2+y^2+2x+8y+17=19 Please show steps :'(

OpenStudy (anonymous):

y do u have a sad face at the end of the ques

OpenStudy (melissa_something):

This is sooo difficult @The.Great.Don

OpenStudy (melissa_something):

What equation is this?

OpenStudy (rock_mit182):

you're setting up oranges with oranges and apples with apples

OpenStudy (rock_mit182):

[(x^2 + 2x + c^2) + (y^2 + 8y + d^2) \]

OpenStudy (rock_mit182):

c^2 = 1 and d^2 = (4)^2

OpenStudy (rock_mit182):

(x^2 + 2x +(1)^2) + (y^2 + 8y + (4)^2)= 19

OpenStudy (rock_mit182):

\[(x^2+ 2x +1) + (y^2 + 8y + 16)=19 \]

OpenStudy (melissa_something):

Yes combining like terms!!

OpenStudy (rock_mit182):

Are you awere of : c^2 + d^2 =17 or (1)^2 + (4)^2 = 17

OpenStudy (rock_mit182):

so we do not need 17 anymore cause we already have it in the equatiaon as the sum of c^2 +d ^2

OpenStudy (rock_mit182):

yeah right

OpenStudy (rock_mit182):

exactly is the clue combining then find the c^2 and b^2 and after that you're almost done

OpenStudy (rock_mit182):

\[(x^2+ 2x + 1^2) \] can be written as: \[(x+1)^2 \] and \[y^2+8y+(4)^2\] can be witten as: \[(y+4)^2 \]

OpenStudy (rock_mit182):

therefore : \[(x+1)^2 + (y+4)^2 = 19 \]

OpenStudy (rock_mit182):

and that's all

OpenStudy (melissa_something):

Wow thank you..

OpenStudy (rock_mit182):

now could you find the center and radius: r ^2 = 19 --> r =sqroot (19)

OpenStudy (rock_mit182):

and the center is way easy..

OpenStudy (melissa_something):

Oh gosh @rock_mit182 im still blurry how you got the 1..& OMG! My bad 19 was supposed to be 9!

OpenStudy (rock_mit182):

did you find the center ?

OpenStudy (rock_mit182):